OpenStudy (anonymous):

The volume of a rectangular box is given by the formula v=LWH. The volume of the box is 120Cu. units. find the width and length. ( L=x+2, width=x, ht=5

OpenStudy (anonymous):

v=LWH 120=(x+2)(x)(5) 120=5x^2+10x 5x^2+10x-120=0 5(x-4)(x+6) 5(4) x=20 w=20 l=22 h=25